Problem Set 11 October 11, 2019 CHEM 1252 4. One of synthetic chemistrys holy grails is to transform methane, CHe(g), into
Answer #1

4.

(a) Below is the balanced reaction of methane reacting with oxygen to release carbon dioxide and water and energy.

CH4 (g) + 2O2 (g) ---->    CO2 (g) + 2H2O (g) + energy

In the above reaction 1 mole of methane in its gaseous form reacts with 2 moles of gaseous oxygen to release 1 mole of carbon dioxide and 2 moles of water molecules as steam and energy,

(b) ΔG° for the above reaction of combustion of methane with oxygen to give carbon dioxide and steam will be
    ΔG° = ΔH - TΔS
Since this reaction is exothermic, ΔH will have a negative value.
For T = 298K
    ΔG° = -802.3 - (298 x -0.005) = -800.8 kJ/mol.

Also as we increase the temperature, ΔG° becomes less negative or we can say ΔG° increases with increase in temperature.